OpenStudy (anonymous):

(15x)+(12y)=292.50 (45x)+(10y)=747.50 Find x and y for both equations. x is the same in both equations y is the same in both equations

OpenStudy (lalipapa123):

(15x)+(12y)=292.50...................(1) (45x)+(10y)=747.50....................(2) ......................... let us use elimination method here divide the first equa. with 45 and the second with 15. so we get 675x+540y=13162.5 675x+150y=11212.5 ....................................... 690y=24375 y=24375/690 y=35.3 substitute the value of y in equ. 1 => (15x)+(12y)=292.50 => 15x+12(35.3)=292.50 => 15x +387.6=292.50 => 15x=387.6-292.50 => 15x =95.1 => x=95.1/15 => x=6.34